# Multi-centre normative brain mapping of intracranial EEG lifespan patterns in the human brain

### v1

@@ 3.4 The effect of age on band power was spatially uniform and frequency band dependent
With the focus shifted to the age model alone, our goal was to understand how age impacted RBP($\cdot$). Taking the most densely populated ROI (the middle temporal region), the age model was implemented for RBP($\cdot$) in each frequency band and regression coefficients for age (denoted $\hat{b}_{age}$) were extracted. A visualisation of model fit in this region is provided in Figure 3A, where blue coloured lines indicate that RBP($\cdot$) decreased with age, whilst red lines indicate an increase with age. This method can be extended to all regions, extracting $\hat{b}_{age}$ for each frequency band and region. The $\hat{b}_{age}$ values can then be visualised on the brain using a hot/cold colour scale, indicating increasing/decreasing RBP($\cdot$) with age (Figure 3B). Finally, disregarding regional information, at the whole-brain level, the age model was implemented in each frequency band and regression coefficients were extracted, along with their 95% confidence intervals (Table 2). Looking at both whole-brain and ROI-level results allowed us to determine if the results had any spatial variation.
Opposing relationships between RBP($\cdot$) and age were observed at both the region-specific and whole-brain levels, with decreasing trends in $\delta$ and $\theta$ frequencies, and increasing trends in the $\alpha$, $\beta$ and $\gamma$ bands.
@@
On the whole-brain level, 95% confidence intervals on $\hat{b}_{age}$ reinforced previous results. Table 2 provides weak-to-no evidence for an age-RBP($\gamma$) relationship and $\theta$’s confidence interval is relatively near zero compared with the remaining three bands. These confidence intervals support the existence of a negative relationship between RBP($\delta$) and age, as well as confidence in a positive one between RBP($\alpha$) and RBP($\beta$), and age. Hence, the effect of age on RBP($\cdot$) was undoubtedly frequency band-specific in this cohort.
The subset of subjects varies between ROIs in the regional analysis; however, the age distribution in each region did not drive any differences between them (Supplementary S4). The trends in Figure 3B persist in a finer-grained parcellation, and when using the original data in which hemispheres were not mirrored (Supplementary S2 and S3).
Figure 3: A) Visualisations of the age model in each frequency band in the most densely populated region, the middle temporal. Blue regression lines indicate a negative value for $\hat{b}_{age}$, whilst red regression lines indicate a positive one. B) Values of $\hat{b}_{age}$ from the age model at the region-level. Values are shown for each ROI and each frequency band of interest. The colour scale is symmetric and fixed across frequency bands with blue representing negative regression coefficients and red representing positive ones. Data from symmetric regions were mirrored, i.e. results are reflected across the midline to provide a whole-brain visualisation.
$\delta$ $\theta$ $\alpha$ $\beta$ $\gamma$
$-0.0007$ $-0.0002$ $0.0004$ $0.0003$ $0.0001$
$(-0.0007,-0.0006)$ $(-0.0002,-0.0001)$ $(0.0004,0.0004)$ $(0.0003,0.0004)$ $(0.0000,0.0001)$
Table 2: Values of $\hat{b}_{age}$ along with 95% confidence intervals for each frequency band. Values are rounded to 4 decimal places and were calculated by applying the age model at the whole-brain level.
@@ 3.5 Certainty about age effect was frequency band specific and impacted by sample size
As a side note, extreme values in Figure 3B co-localise to low sampled regions seen in Figure 4, e.g., the frontal pole stands out from the general trend in $\beta$ and only has $n=37$, highlighting the importance of sample size.
@@
Figure 4: Summaries for each ROI under consideration. The $x$-axis displays the region name and is ordered by the number of subjects per region. A) number of subjects. B) a binary indicator of whether the 95% confidence interval for $\hat{b}_{age}$ contains 0 in each frequency band. Singular fits given in black. C) standard errors of $\hat{b}_{age}$ in every frequency band.
@@
With the exception of the $\theta$-band, these findings became robust in ROIS with $n\geq 295$. This suggests, that equal and large sample sizes in all ROIs may lead to more uniform results and provides further confidence in the relationships we have identified between age and $\delta$, $\alpha$ and $\beta$ bands, and the likely lack of one between RBP($\gamma$) and age.
The age model returned a singular fit in a few regions (Figure 4B). These regions typically had a low number of data points and a small number of sites. We retained all models in all ROIs for completeness.
@@ 4 Discussion
In this study, we considered how age, sex, and recording site, might impact RBP($\cdot$) in the setting of a normative icEEG map. To understand the effects of the variables, various LMMs were fitted, and it was concluded that sex and RBP($\cdot$) were not related in this cohort, whilst RBP($\cdot$) values present a moderate relationship with subject age and a notable relationship with the recording site.
@@
This work provides insight into what the trajectory of RBP($\cdot$) extracted from icEEG might look like in healthy lifespan. Additionally, it confirms that site effects must be considered when modelling multi-centre data, whilst also demonstrating that sex perhaps need not be considered when analysing normative icEEG RBP($\cdot$).
4.1 The relationship between site, sex, age and band power
@@
The RAM data set is publicly available and comprises EEG, as well as other data types, from nine sites. Hence, it is regularly used in academic work, producing a range of results (Taylor et al., 2022, Wang et al., 2023, Nozari et al., 2020, Kozma et al., 2023, Goldstein et al., 2019, Das et al., 2022). Whilst the multi-site challenge has been recognised before in Miller (2019), to our knowledge, this and other studies have not statistically accounted for the multi-site structure present in their data. Our work reveals that, in the normative icEEG setting at least, the site effect can account for up to one-quarter of the variation in the response variable, demonstrating the necessity of considering such effects.
Several factors might be driving these site differences, such as recording techniques. Concerning grid and strip versus depth electrodes, each has advantages and disadvantages (Taussig et al., 2015), and past work found that while both were efficient presurgical localisation techniques, seizure freedom occurred at a higher rate following the use of depth electrodes, evidencing a difference between the two (Toth et al., 2019). In our cohort (following all preprocessing), 12 sites involved a mixture of electrode types, while 3 contained only depth electrodes, perhaps driving some of the site variation. Other possibilities include race, geography and subject selection criteria. Previous research found race and marital status to be marginally associated with whether or not an individual proceeds to epilepsy surgery (Berg et al., 2003), so it is plausible such factors could also impact whether individuals undergo intracranial examination.
@@
Moving to the fixed effects, our work produced convincing results in the $\delta$, $\alpha$ and $\beta$ bands when it came to a relationship between RBP($\cdot$) and age. Specifically, RBP($\delta$) decreased with subject age, while RBP($\alpha$) and RBP($\beta$) increased. Generally, we found an increase in RBP($\cdot$) with age in fast bands and a decrease in RBP($\cdot$) with age in slower bands.
@@
These fixed effed results concur with previous literature on scalp EEG, although here the age range is reduced to children and adolescents, which found fast bands increased and the slow bands decreased in RBP($\cdot$) (Gasser et al., 1988, Clarke et al., 2001). Using the same age range, MEG PSD maps found an analogous relationship between power and age but reported sex differences in $\delta$, $\beta$ and $\alpha$ (Ott et al., 2021). They do not report a main effect of sex in $\theta$ and $\gamma$, which are the same bands in which we found no notable relationship for either fixed effect.
@@
Looking at the full lifespan (7-84 years) using MEG, previous work found identical trends (a decrease in RBP($\delta$) and RBP($\theta$), and the converse for higher bands); however, they report slight changes around the sixth decade of life (Gómez et al., 2013). Since our maximum age is 18 years less at 66, this result still complements our findings. Further, a study using a small set of scalp electrodes across a large adult cohort found the same age trajectories as ours between 20-40 years, including the strongest decrease being in the $\delta$-band (Hashemi et al., 2016). However, they found that these effects diminished in older age ranges and also reported significant sex differences, which we did not. Nevertheless, they use a different, although related, feature and modality, which may facilitate the sex differences.
@@ 4.2 Existing normative mapping work and how it relates to the ageing patterns
The majority of normative mapping studies outlined consider their maps from a static viewpoint and did not incorporate potential age effects. In this study, we demonstrated that RBP($\cdot$) varies with age and that this variation is frequency band-specific. This is not surprising, as there is much previous work evidencing that healthy ageing does impact brain activity in modalities such as MEG and scalp EEG (Gomez et al., 2017, Gómez et al., 2013, Duffy et al., 1993). Therefore we suggest, based on our findings, that future work does not consider normative maps as a static snapshot and instead strives to incorporate dynamic features, such as age, into any study.
@@ 4.3 Clinical potential
Turning now to an application of normative maps, the comparison of individuals with epilepsy (or other disorders) to normalised healthy controls is a common one. It is gaining traction in the field of neurophysiology, using modalities which facilitate said controls such as scalp EEG and MEG (Bosch-Bayard et al., 2020a, Owen et al., 2023b, Janiukstyte et al., 2023, Owen et al., 2023a). Similar work has been produced in the invasive icEEG setting; hence, it is plausible that the ageing trends demonstrated here could yield similar clinical potential (Taylor et al., 2022, Bernabei et al., 2022, Kozma et al., 2023).
@@ 4.4 Limitations and future directions
The concept of ‘normative’ data collected from individuals with epilepsy is perhaps contentious. These methods are defensible (Rutherford et al., 2023) and have been applied in previous work (Bernabei et al., 2022, Frauscher et al., 2018a, Kalamangalam et al., 2020, Kozma et al., 2023, Taylor et al., 2022, Wang et al., 2023); however, epilepsy is increasingly defined as a network disorder (Rayner and Tailby, 2017, Kramer and Cash, 2012, Bernhardt et al., 2015), which contrasts with the notion of delineating normal and abnormal tissue. In particular, previous work has shown that more complete resection of seizure onset regions is not associated with more favourable surgical outcomes (Gascoigne et al., 2023). Hence, it is arguable that the icEEG description of pathological tissue is complex and incomplete.
Sample size impacted results despite our efforts in data collection (Section 3.5), with high regional variation in number of subjects. In an ideal world, we would attain a similar number of subjects per region so they could be more accurately compared. In practice, however, this is not feasible due to some areas being more prone to pathology and therefore being more likely to have electrodes implanted – for example, drug-resistant epilepsies in adults are common in the temporal lobe (Bernhardt et al., 2019).
@@
Combining the discussions around frequency ranges and mode complexity, there is also evidence that more granular frequency bands could affect the model selection process. For example, using healthy MEG data to determine how RBP($\cdot$) changes with age, Gómez et al. (2013) found that a linear regression model performed best for RBP(low-$\beta$), but a quadratic model is preferable for RBP(high-$\beta$). Analogous results were found in scalp EEG in the $\alpha$-band as well as the $\beta$-band (more complex models were required for the upper end of the band range) (Gasser et al., 1988).
Different features of EEG could also be considered, in particular, absolute band power as an alternative to the relative band power used here. A previous EEG study found that both relative and absolute $\delta$-power decreased with age, but absolute power failed to yield significant results in other bands, whilst relative power did not (Clarke et al., 2001). Further, one study found RBP($\cdot$) is the optimal choice for diagnosis using EEG, perhaps suggesting absolute band power need not be investigated (Psatta et al., 2000).

### v2

@@ 3.4 The effect of age on band power was spatially uniform and frequency band dependent
With the focus shifted to the age model alone, our goal was to understand how age impacted RBP($\cdot$).
Taking the most densely populated ROI (the middle temporal region), the age model was implemented for RBP($\cdot$) in each frequency band and regression coefficients for age (denoted $\hat{b}_{age}$) were extracted. A visualisation of model fit in this region is provided in Figure 5A, where blue coloured lines indicate that RBP($\cdot$) decreased with age, red lines indicate an increase with age and black lines indicate no relationship (based on 95% confidence intervals for $\hat{b}_{age}$). This method can be extended to all regions, extracting $\hat{b}_{age}$ for each frequency band and region. The $\hat{b}_{age}$ values can then be visualised on the brain using a hot/cold colour scale, indicating increasing/decreasing RBP($\cdot$) with age (Figure 5B). Finally, disregarding regional information, at the whole-brain level, the age model was implemented in each frequency band and regression coefficients were extracted, along with their 95% confidence intervals (Table 2). Looking at both whole-brain and ROI-level results allowed us to determine if the results had any spatial variation.
Figure 5: A) Visualisations of the age model in each frequency band in the most densely populated region, the middle temporal. Using the 95% confidence intervals for $\hat{b}_{age}$ in the region, blue regression lines indicate a negative relationship between RBP($\cdot$) and age, whilst red regression lines indicate a positive one and black indicates no relationship. B) Values of $\hat{b}_{age}$ from the age model at the region-level. Values are shown for each ROI and each frequency band of interest. The colour scale is symmetric and fixed across frequency bands with blue representing negative regression coefficients and red representing positive ones. Data from symmetric regions were mirrored, i.e. results are reflected across the midline to provide a whole-brain visualisation.
$\delta$ $\theta$ $\alpha$ $\beta$ $\gamma$
$-6.1\times 10^{4}$ $-0.8\times 10^{4}$ $4.1\times 10^{4}$ $2.9\times 10^{4}$ $-0.2\times 10^{4}$
$(-6.7,-5.4)\times 10^{4}$ $(-1.2,-0.4)\times 10^{4}$ $(3.8,4.5)\times 10^{4}$ $(2.4,3.3)\times 10^{4}$ $(-0.7,0.3)\times 10^{4}$
Table 2: Values of $\hat{b}_{age}$ along with 95% confidence intervals for each frequency band. Values are rounded to 5 decimal places and were calculated by applying the age model at the whole-brain level.
Opposing relationships between RBP($\cdot$) and age were observed at both the region-specific and whole-brain levels, with decreasing trends in $\delta$ and $\theta$ frequencies, increasing trends in the $\alpha$ and $\beta$ bands and weaker results found for the $\gamma$ band.
@@
On the whole-brain level, 95% confidence intervals on $\hat{b}_{age}$ reinforced previous results. Table 2 provides no evidence for an age-RBP($\gamma$) relationship and $\theta$’s confidence interval is relatively near zero compared with the remaining three bands. These confidence intervals support the existence of a negative relationship between RBP($\delta$) and age, along with a positive one between RBP($\alpha$) and RBP($\beta$), and age. Hence, the effect of age on RBP($\cdot$) was undoubtedly frequency band-specific in this cohort.
The subset of subjects varies between ROIs in the regional analysis; however, the age distribution in each region did not drive any differences between them (Supplementary S9). The trends in Figure 5B persist in a finer-grained parcellation, and when using the original data in which hemispheres were not mirrored (Supplementary S3 and S4).
@@ 3.5 Certainty about age effect was frequency band specific and impacted by sample size
As a side note, extreme values in Figure 5B co-localise to low sampled regions seen in Figure 6, e.g., the frontal pole stands out from the general trend in $\theta$ and only has $n=38$, highlighting the importance of sample size.
@@
With the exception of the $\theta$-band, these findings became robust in ROIs with $n\geq 291$. This suggests that equal and large sample sizes in all ROIs may lead to more uniform results and provides further confidence in the relationships we have identified between age and $\delta$, $\alpha$ and $\beta$ bands, and the lack of one between RBP($\gamma$) and age.
The age model returned some singular fits at the regional level (Figure 6B). These regions typically had a low number of data points and a small number of hospitals. We retained all models in all ROIs for completeness.
Figure 6: Summaries for each ROI under consideration. The $x$-axis displays the region name and is ordered by the number of subjects per region. A) number of subjects. B) a binary indicator of whether the 95% confidence interval for $\hat{b}_{age}$ contains 0 in each frequency band. Singular fits given in black. C) standard errors of $\hat{b}_{age}$ in each frequency band.
@@ 4 Discussion
In this study, we considered how age, sex, and recording hospital, might impact RBP($\cdot$) in the setting of a normative icEEG map. To understand the effects of the variables, various LMMs were fitted, and it was concluded that sex and RBP($\cdot$) were not related in this cohort, whilst RBP($\cdot$) values present a moderate relationship with subject age and a notable relationship with the recording hospital.
@@
This work provides insight into what the trajectory of RBP($\cdot$) extracted from icEEG might look like in healthy lifespan. Additionally, it confirms that hospital site effects must be considered when modelling multi-centre data, whilst also demonstrating that sex perhaps need not be considered when analysing normative icEEG RBP($\cdot$).
4.1 The relationship between hospital, sex, age and band power
Hospital and band power
In other neuroscience research concerned with normative mapping, such as neuroimaging, the multi-centre problem is widely recognised (Hu et al., 2023, Keshavan et al., 2016, Jovicich et al., 2006). It is common practice to model hospital site effects (or ‘batch’ effects in neuroimaging literature) as random offsets in normative models (Little et al., 2024, Hibar et al., 2018, Ge et al., 2024, Bethlehem et al., 2022, Kim et al., 2022). Whilst it is impossible to capture all sources of hospital (or scanner) heterogeneity, there is some consideration of them in this field. Meanwhile, other neuroscience studies, such as those in EEG, do not typically consider hospital site effects.
@@ 4.1 The relationship between hospital, sex, age and band power
For example, the RAM data set involves nine hospitals and is regularly used in academic work to produce results in icEEG (Taylor et al., 2022, Wang et al., 2023, Nozari et al., 2023, Kozma et al., 2024, Goldstein et al., 2019, Das et al., 2022). Whilst the multi-centre challenge here has been recognised before by Miller (2019), to our knowledge, this and other studies have not statistically accounted for the multi-centre structure present. Our work reveals that, in the relative band power properties, the size of the hospital effect is frequency band specific, accounting for up to 30% of the variation.
This demonstrates the necessity of hospital effect consideration, but also raises the question of what drives the differences. Supplementary S8 considers the metadata we have available – electrode type, age range and originating cohort (RAM/other) – as possible contributors to recording hospital differences, but finds no systematic variation in these factors. Other technical aspects could be involved, such as recording settings or electrode manufacturer. Demographic possibilities include race, geography and the hospital’s subject selection criteria.
Whilst both depth and grid/strip electrodes have differing advantages (Taussig et al., 2015), we found no evidence that electrode type was driving recording hospital variation (Supplementary S8). Past work found that both types were efficient for presurgical evaluation, and Frauscher et al. (2018a) do not find any change in spectral density due to electrode type.
Our data did not contain information on subject ethnicity, however, this data should be collected in the future, to address the issue of past exclusionary practices in neuroscience (Ricard et al., 2022, Li et al., 2022). Previous research found race and marital status to be marginally associated with whether or not an individual proceeds to epilepsy surgery (Berg et al., 2003), so it is plausible such factors could also impact whether individuals undergo intracranial examination.
Finally, we wish to highlight the dependency of the hospital site effect size on signal feature in our cohort (Table 1). For example, the hospital site effect was weakest in RBP($\alpha$). Investigating whether this is reproducible in other data sets is a further avenue of future research. There may also exist signal features which are more robust towards the effect of recording hospital.
In summary, although the multi-centre effect is under-explored in icEEG research, it is present in our work. Future work should improve on our own by collecting more metadata, to help identify what is driving the hospital site effect and to explain why different signal properties yield different results.
Age and band power
@@
Moving to the fixed effects, our work produced convincing results in the $\delta$, $\alpha$ and $\beta$ bands when testing for a relationship between RBP($\cdot$) and age. Specifically, RBP($\delta$) decreased with subject age, while RBP($\alpha$) and RBP($\beta$) increased. More generally, we found an increase in RBP($\cdot$) with age in fast bands and a decrease in RBP($\cdot$) with age in slower bands.
@@
These fixed effed results concur with previous literature on scalp EEG, although here the age range is reduced to children and adolescents, which found fast bands increased and the slow bands decreased in RBP($\cdot$) (Gasser et al., 1988, Clarke et al., 2001). Using the same age range, MEG power spectral density maps found an analogous relationship between power and age (Ott et al., 2021).
@@
Looking at the full lifespan (7-84 years) using MEG, previous work found identical trends (a decrease in RBP($\delta$) and RBP($\theta$), and the converse for higher bands); however, the authors report slight changes around the sixth decade of life (Gómez et al., 2013). Since our maximum age is 18 years less at 66, this result still complements our findings. Further, a study using a small set of scalp electrodes across a large adult cohort found the same age trajectories as ours between 20-40 years, including the strongest decrease being in the $\delta$-band (Hashemi et al., 2016). However, they found that these effects diminished in older age ranges, which our choice of linear modelling does not facilitate.
Finally, there is evidence in the literature of a decrease in alpha band power with age – Tröndle et al. (2023) and Turner et al. (2023) report this across adulthood, and Whitford et al. (2007) in adolescence. However, these studies use absolute power rather than relative, in scalp rather than intracranial EEG, so this does not contradict our findings. Tröndle et al. (2023) also argue that the decrease has been overestimated due to a lack of consideration of the aperiodic component.
Sex and band power
Our work did not find any evidence for a sex-RBP($\cdot$) relationship in this cohort. In the literature however, Ott et al. (2021) report sex differences in MEG power spectral density maps for children and adolescents in $\delta$, $\beta$ and $\alpha$. They do not report a main effect of sex in $\theta$ and $\gamma$, which are the same bands in which we found no notable relationship for either fixed effect. Further, Hashemi et al. (2016) report significant sex differences in total band power with age in an adult cohort using scalp EEG.
Nevertheless, as with the evidence for the lifespan decrease in alpha power, both studies use a different, although related, feature and modality, which may facilitate the sex differences. Finding a difference in absolute power (Hashemi et al., 2016), but not relative power (our cohort) may suggest that the distribution of power is the same for men and women, but that the values differ. Alternatively, it is plausible that there is a genuine effect of sex on RBP($\cdot$) but that it is too small to detect in our cohort.
@@ 4.2 Existing normative mapping work and how it relates to the ageing patterns
The majority of normative mapping studies outlined consider their maps from a static viewpoint and did not incorporate potential age effects. In this study, we demonstrated that RBP($\cdot$) varies with age and that this variation is frequency band-specific. This is not surprising, as there is much previous work evidencing that healthy ageing does impact brain activity in modalities such as MEG and scalp EEG (Gomez et al., 2017, Gómez et al., 2013, Duffy et al., 1993). Therefore we suggest, based on our findings, that future work no longer considers icEEG normative maps as a static snapshot and instead strives to incorporate dynamic features, such as age, into any study.
@@ 4.3 Clinical potential
Turning now to an application of normative maps, the comparison of individuals with epilepsy (or other disorders) to normalised healthy controls is a common one. It is gaining traction in the field of neurophysiology, using modalities which facilitate said controls such as scalp EEG and MEG (Bosch-Bayard et al., 2020a, Owen et al., 2023b, Janiukstyte et al., 2023, Owen et al., 2023a). Similar work has been produced in the invasive icEEG setting; hence, it is plausible that the ageing trends demonstrated here could yield similar clinical potential (Taylor et al., 2022, Bernabei et al., 2022, Kozma et al., 2024).
@@ 4.4 Limitations and future directions
The concept of ‘normative’ data collected from individuals with epilepsy is perhaps contentious. These methods are defensible (Rutherford et al., 2023) and have been applied in previous work (Bernabei et al., 2022, Frauscher et al., 2018a, Kalamangalam et al., 2020, Kozma et al., 2024, Taylor et al., 2022, Wang et al., 2023); however, epilepsy is increasingly defined as a network disorder (Rayner and Tailby, 2017, Kramer and Cash, 2012, Bernhardt et al., 2015), which contrasts with the notion of delineating normal and abnormal tissue. In particular, previous work has shown that more complete resection of seizure onset regions is not associated with more favourable surgical outcomes (Gascoigne et al., 2024). Hence, it is arguable that the icEEG description of pathological tissue is complex and incomplete.
Additionally, sample size impacted results despite our efforts in data collection (Section 3.5), with high regional variation in number of subjects. In an ideal world, we would attain a similar number of subjects per region so they could be more accurately compared. In practice, however, this is not feasible due to some areas being more prone to pathology and therefore being more likely to have electrodes implanted – for example, drug-resistant epilepsies in adults are common in the temporal lobe (Bernhardt et al., 2019).
@@
Combining the discussions around frequency ranges and model complexity, there is also evidence that more granular frequency bands could affect the model selection process. For example, using healthy MEG data to determine how RBP($\cdot$) changes with age, Gómez et al. (2013) found that a linear regression model performed best for RBP(low-$\beta$), but a quadratic model is preferable for RBP(high-$\beta$). Analogous results were found in scalp EEG in the $\alpha$-band as well as the $\beta$-band (more complex models were required for the upper end of the band range) (Gasser et al., 1988).
Different features of EEG could also be considered, in particular, absolute band power as an alternative to the relative band power used here. This would allow for direct comparison to the aforementioned work which finds sex differences in absolute power (Hashemi et al., 2016). However, a previous EEG study investigating age, sex and band power found that absolute power failed to find significant lifespan effects in all bands except $\delta$, whilst relative power succeeded, suggesting further investigation using absolute power is unnecessary (Clarke et al., 2001).